# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ 4.1 General bifurcation results
The proof proceeds via the equivariant branching lemma, see for example [CL00, Theorem 2.3.2]. By Equation 29, $B(u)$ is elliptic, hence the linearization of $G(u,\Lambda)$ is elliptic as well, and hence the upper left $2\times 2$ block of ${\partial}_{U}{\cal G}$ is a Fredholm operator. At every $\lambda_{2}$ from (51), the zero eigenvalue is of finite multiplicity, namely multiplicity 2 in (a,b), and multiplicity 4 in (c), and the symmetries of the bifurcating branches are as follows:
  1. (a)
    $X(\varepsilon)$ admits a new non trivial symmetry wrt translations in $x$. Assume that $A_{-m}=A_{m}$ with $A_{m}\in{\mathbb{R}}$, so that $u=\cos(k_{m}x)$. The (not normalized) normal vector is
    $\nu={\partial}_{x}X(\varepsilon)\times{\partial}_{\varphi}X(\varepsilon)$ $=\begin{pmatrix}1\\ -k_{m}\sin(k_{m}x)\cos(\varphi)\\ -k_{m}\sin(k_{m}x)\sin(\varphi)\end{pmatrix}\times\begin{pmatrix}0\\ -(1+\cos(k_{m}x))\sin(\varphi)\\ (1+\cos(k_{m}x))\sin(\varphi)\end{pmatrix}$
    $=-(1+\cos(k_{m}x))\begin{pmatrix}k_{m}\sin(k_{m}x)\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}.$
    Hence $\left\langle\nu,e_{j}\right\rangle\neq 0$ for $j=1,2,3$, hence $\rm{span}\{\left\langle\nu,e_{j}\right\rangle\}$ is a subset of the generator of the center manifold of $X(\varepsilon)$. Following this with
    $X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\cos(x))\sin(\varphi)\\ (1+\cos(x))\cos(\varphi)\end{pmatrix},$
    we have that $\left\langle X(\varepsilon)\times e_{1},\nu\right\rangle=0$.
  2. (b)
    As for (a) we can calculate the center manifold in the bifurcating branch, with
    $\nu={\partial}_{x}X(\varepsilon)\times{\partial}_{\varphi}X(\varepsilon)=\begin{pmatrix}0\\ 2\sin(2\varphi)\sin(\varphi)+\cos(2\varphi)\cos(\varphi)\\ 2\sin(2\varphi)\cos(\varphi)+\cos(2\varphi)\sin(\varphi)\end{pmatrix}$
    and
    $\left\langle X(\varepsilon)\times e_{1},\nu\right\rangle=\left\langle\begin{pmatrix}0\\ (1+\cos(2\varphi))\sin(\varphi)\\ -(1+\cos(2\varphi))\cos(\varphi)\end{pmatrix},\nu\right\rangle=\cos(2\varphi)(1+\cos(2\varphi))\cos(\varphi)\sin(\varphi).$
    Hence in the center manifold of $X(\varepsilon)$ are besides the translation in $y$ and $z$ direction also rotations around the $x$ axis.
  3. (c)
    The 4D kernel generates two different types of branches. For the coils we have
    ${\partial}_{x}X(\varepsilon)$ $=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}-k_{m}\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}$
    ${\partial}_{\varphi}X(\varepsilon)$ $=(1+\cos(k_{m}x+n\varphi)\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}-n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}$
@@
    $\nu=$ $(1+\cos(k_{m}x+n\varphi)\begin{pmatrix}0\\ -\cos(\varphi)\\ -\sin(\varphi)\end{pmatrix}-n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}$
    $-k_{m}\sin(k_{m}x+n\varphi)(1+\cos(k_{m}x+n\varphi))\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$
    Obviously, this results in a non trivial translational symmetry in $x$. Similarly
    $X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\cos(k_{m}x+n\varphi))\sin(\varphi)\\ (1+\cos(k_{m}x+n\varphi))\cos(\varphi)\end{pmatrix},$
    such that $\left\langle X(\varepsilon)\times e_{1},\nu\right\rangle=n\sin(k_{m}x+n\varphi)(1+\cos(k_{m}x+n\varphi))$ yields the same action as the translation. For the buckling solutions similar computations show that $\mathrm{span}\{\left\langle\nu,e_{1}\right\rangle,\left\langle X(\varepsilon)\times e_{1},\nu\right\rangle\}$ is two dimensional, hence the full symmetry is needed.
@@ 4.2 Reduced equation
To further describe the bifurcations from Theorem 4.1, we derive the reduced equations (or amplitude equations AEs) at ${\cal C}_{L}$. These are straight forward computations, but somewhat cumbersome in our geometric setting and thus we partly use Maple. Here we explain the basic steps, and the computations for the pearling and wrinkling instability, but for the coils vs buckling we relegate some details to Appendix LABEL:aeapp. Additionally we give the numerical values of the coefficients of the AEs for some of examples, which will later be compared to numerics for the full problem.
According to Theorem 4.1, at “simple” (strictly speaking of multiplicity two, but from here we factor out the translation or rotation) bifurcation points (51), i.e., $(m,n)=(m,0)$ or $(m,n)=(0,n)$, the bifurcating steady state branch is tangential to solutions of $D{\cal G}(0,\Lambda)[(\psi,0)]=0$ where ${\cal G}$ is the rhs of (49). We thus introduce the small parameter
@@
$u=\varepsilon A\psi_{m,n}+\varepsilon^{2}\sum_{(i,j)\in{\cal B}}A_{i,j}\psi_{i,j}+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),$ (57)
@@
Near a bifurcation point from a stable branch, the critical mode determines the evolution, while all others are exponentially damped. Hence, in a standard AE setting, the signs of $a$, $\beta_{2}$ and $b$ in (59) determine the stability on the bifurcating branch. Here, from (59) we get stability information for steady states under area constraint, with $\lambda_{2}$ as extrinsic fixed parameter, but this can be extended to stability under area and volume constraints as follows. Close to the bifurcation point $\lambda_{2}(m,n)$ the solution reads $X(\varepsilon)={\cal C}_{L}-\varepsilon A\psi N$. Then the solution is linearly stable iff ${\partial}_{U}{\cal G}(X(\varepsilon),\Lambda)$ (which includes the $q_{1}$ area constraint) has only negative eigenvalues. For admissible (area and volume preserving) perturbations $\phi:{\cal C}_{L}\rightarrow{\mathbb{R}}$ of $X(\varepsilon)$, we have the necessary and sufficient conditions
$\int_{X(\varepsilon)}H(\varepsilon A\psi)\phi\,{\rm d}S=\int_{X(\varepsilon)}\varepsilon A\psi g^{-1/2}(\varepsilon A\psi)\phi\,{\rm d}S=0.$
For the mean curvature we have $H(\varepsilon A\psi)=\frac{1}{2}-\varepsilon A(\mu(m,n)+1)\psi+{\cal O}(\varepsilon^{2})$, and hence admissible perturbations must be orthogonal to $\psi$. Thus, near bifurcation points associated with loss of stability of ${\cal C}_{L}$ we have the following results:
  • Branches bifurcating from simple bifurcation points (pearling or wrinkling), are stable wrt (7).
  • Solutions on branches from a double bifurcation point (buckling and coiling), can only be destabilized by the other kernel vectors. We will see examples of this in §5.
@@ 4.2.1 Pearling instability
[…]
@@
$b=$ $\frac{5}{\left(k^{2}{-}1\right)k^{2}\left(4k^{4}{-}5k^{2}{+}4c_{0}{-}1\right)}\Bigg{(}\frac{3}{20}{+}k^{14}{+}\left(4c_{0}{-}\frac{73}{20}\right)k^{12}{+}\left(\frac{8c_{0}}{5}{-}\frac{13}{10}\right)k^{10}{+}\left(\frac{32}{5}c_{0}^{2}{-}\frac{52}{5}c_{0}{+}\frac{73}{20}\right)k^{8}$
${+}\left(\frac{16}{5}c_{0}^{2}{-}8c_{0}{+}\frac{31}{5}\right)k^{6}{+}\left(\frac{4c_{0}}{5}{-}\frac{43}{20}\right)k^{4}{+}\left(\frac{12c_{0}}{5}{-}\frac{3}{2}\right)k^{2}\Bigg{)}.$
Recall that here $k=k_{m}=2m\pi/L$, hence $a<0$ for $L>2m\pi$, and similarly $b<0$ for non small $L$. For instance, $a\approx-0.6$ and $b\approx-3$ in Fig.5, where we compare the AE predictions for $m=1$ pearling with $c_{0}=0$ and $L=10$ with numerics. In any case for $ab>0$ we set $\beta_{2}=-1$ (subcritical case), and otherwise $\beta_{2}=1$ (supercritical case), giving the amplitude $A=\sqrt{\left|\frac{a}{b}\right|}$, with also $\alpha_{2}$ uniquely determined. The change of the volume can be approximated by inserting the amplitude Ansatz, giving $0=A_{0}+A^{2}+\beta_{2}$ at $\varepsilon^{2}\psi_{00}$, and comparison with (60) shows that preserving the area and the volume is not possible along the pearling branch.
@@ 4.2.2 Wrinkling instability
[…]
@@ 4.2.3 Coiling/buckling instability
For the instability with critical wave vector $(m,n)=(1,1)$ and $(m,n)=(1,-1)$ we choose the ansatz
@@
[…]
From that we can directly recover the two types of solutions: Buckling with $A=\sqrt{\left|\frac{a}{b_{1}}\right|}$ and $B=0$ or the other way round; coiling with $A=B=\pm\sqrt{\left|\frac{a}{b_{1}+b_{2}}\right|}$. Here again $\beta_{2}=\pm 1$, depending on the signs of $a$, $b_{1}$, and $b_{1}+b_{2}$.
We have $a=k_{m}^{2}$, but the computation of the coefficients $b_{1}$ and $b_{2}$ naturally is more cumbersome than for the scalar case (pearling and wrinkling), and therefore relegated to Appendix LABEL:aeapp. Here we instead briefly comment on the bifurcation directions, and stability of the bifurcating branches under volume and area preserving flows, see also Fig.3. The Jacobian of the rhs of (64) at a steady state $(A,B)$ is
@@
For $a>0$, on the coil solution branch the eigenvalues are $\sigma_{1}=(a\beta_{2}+3{\rm sign}(b_{1})|a\beta_{2}|)$ and $\sigma_{2}=(a\beta_{2}+\frac{b_{2}}{|b_{1}|}|a\beta_{2}|)$. Hence the coils are stable for $\frac{b_{2}}{|b_{1}|}<-{\rm sign}(\beta_{2})$.
For the buckling the Jacobian has the eigenvalues
$\sigma_{1}=2\frac{b_{2}+b_{1}}{|b_{1}+b_{2}|}\quad\sigma_{2}=\frac{b_{2}-b_{1}}{|b_{1}+b_{2}|},$
where $\sigma_{1}$ corresponds to the eigenvector $v_{1}=(1,1)^{T}$ and $\sigma_{2}$ to $v_{2}=(1,-1)^{T}$. Hence the relevant eigenvalue is $\sigma_{2}$, which is negative for $b_{2}<b_{1}$. In §5 we give two comparisons of the predictions from (64) with numerics, namely $c_{0}$ and $L=10$ yielding $(a,b_{1},b_{2})\approx(0.39,-0.67,-0.82)$ in Fig.4, and $c_{0}=0.48$ and $L=10$ yielding $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$ in Fig.9. In both cases we get the correct branching and stability predictions, and also very good quantitative agreement.
@@
Figure 3: Schematic linearized stability diagram of the amplitude equation depending on $b_{1}$ and $b_{2}$ Orange stands for the coil and green for the wave branch. Thick lines indicate stable solutions.

### v2

@@ 4.1 General bifurcation results
The proof proceeds via the equivariant branching lemma; see, for example, [CL00, Theorem 2.3.2]. By (29), $B(u)$ is elliptic; hence the linearization of $G(u,\Lambda)$ is elliptic as well, and thus the upper left $2\times 2$ block of $\partial_{U}{\cal G}$ is a Fredholm operator. At every $\lambda_{2}$ from (53), the zero eigenvalue is of finite multiplicity, namely multiplicity 2 in (a,b), and multiplicity 4 in (c), with the symmetries acting on the amplitudes $(A_{mn},A_{m-n})\in{\mathbb{C}}^{2}$. For (c), the joint reflection gives the condition that either $A_{mn}$ or $A_{m-n}$ must be zero (coiling), while enforcing a $Z_{2}\times Z_{2}$ symmetry yields $A_{mn}=A_{m-n}$ (buckling). In all cases, the isotropy subgroups acting on the amplitudes yield one-dimensional ${\rm Fix}(\Sigma)$. ∎
Remark 4.2.
The equivariant group actions correspond to the rigid body motions stated in (42). As the solution $X$ on the bifurcating branch has the same symmetry group as ${\cal C}-uN$, with $u$ one of the functions from Theorem 57, we shall compute that, for example, in (a), the translations in $x$ are nontrivial rigid body motions of $X-(\varepsilon A_{m}e^{ik_{m}x}+\mathrm{c.c.}+{\cal O}(\varepsilon^{2}))N$.
Assume that $A_{-m}=A_{m}$ with $A_{m}\in{\mathbb{R}}$, so that $u=\cos(k_{m}x)$. The (not normalized) normal vector is
$\nu=\partial_{x}X(\varepsilon)\times\partial_{\varphi}X(\varepsilon)$ $=\begin{pmatrix}1\\ -\varepsilon k_{m}\sin(k_{m}x)\cos(\varphi)\\ -\varepsilon k_{m}\sin(k_{m}x)\sin(\varphi)\end{pmatrix}\times\begin{pmatrix}0\\ -(1+\varepsilon\cos(k_{m}x))\sin(\varphi)\\ (1+\varepsilon\cos(k_{m}x))\cos(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2})$
$=-(1+\varepsilon\cos(k_{m}x))\begin{pmatrix}\varepsilon k_{m}\sin(k_{m}x)\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}).$
Hence $\langle\nu,e_{j}\rangle\neq 0$ for $j=1,2,3$; thus, ${\rm span}\{\langle\nu,e_{j}\rangle,j=1,2,3\}$ is a subspace of the tangent space of the center manifold of $X$. However, as this coincides with the number of equivariant group actions close to the bifurcation point, the tangent space coincides with the three translations. In contrast, a rotation around the $x$-axis yields
$X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\varepsilon\cos(k_{m}x))\sin(\varphi)+{\cal O}(\varepsilon^{2})\\ (1+\varepsilon\cos(k_{m}x))\cos(\varphi)+{\cal O}(\varepsilon^{2})\end{pmatrix}$,
so $\langle X(\varepsilon)\times e_{1},\nu\rangle=0$ and the rigid body motion is trivial. A similar argument applies for (b).
In (c), the 4D kernel generates two different types of branches. For the buckling, a combination of the above arguments yields that translations in $x$ and the rotations in $\varphi$ are nontrivial rigid body motions. For the coiling we have, choosing $u=\cos(k_{m}x+n\varphi)$,
$\partial_{x}X(\varepsilon)$ $=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}-\varepsilon k_{m}\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
$\partial_{\varphi}X(\varepsilon)$ $=(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}-\varepsilon n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ \cos(\varphi)\\ \sin(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
@@
$\nu=$ $(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}0\\ -\cos(\varphi)\\ -\sin(\varphi)\end{pmatrix}-\varepsilon n\sin(k_{m}x+n\varphi)\begin{pmatrix}0\\ -\sin(\varphi)\\ \cos(\varphi)\end{pmatrix}$
$-\varepsilon k_{m}\sin(k_{m}x+n\varphi)(1+\varepsilon\cos(k_{m}x+n\varphi))\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}+{\cal O}(\varepsilon^{2}).$
This results in a nontrivial translational symmetry in $x$. Similarly,
$X(\varepsilon)\times e_{1}=\begin{pmatrix}0\\ -(1+\cos(k_{m}x+n\varphi))\sin(\varphi)\\ (1+\cos(k_{m}x+n\varphi))\cos(\varphi)\end{pmatrix}+{\cal O}(\varepsilon^{2}),$
such that $\langle X\times e_{1},\nu\rangle=\varepsilon n\sin(k_{m}x+n\varphi)(1+\varepsilon\cos(k_{m}x+n\varphi))+{\cal O}(\varepsilon^{2})$ yields the same action as the translation. These computations are also important for the numerics, as they give the formulas and numbers of phase conditions needed for the numerical continuation in §5. $\rfloor$
@@ 4.2 Reduced equation
To further describe the bifurcations from Theorem 4.1, we derive the reduced equations (or amplitude equations AEs) at ${\cal C}_{L}$. These are straight forward computations, but somewhat cumbersome in our geometric setting and thus we partly use Maple. Here we explain the basic steps, and the computations for the pearling and wrinkling instability, but for the coiling vs buckling we relegate some details to Appendix A. Additionally we give the numerical values of the coefficients of the AEs for some of examples, which will later be compared to numerics for the full problem.
According to Theorem 4.1, at “simple” (strictly speaking of multiplicity two, but from here we factor out the translation or rotation) bifurcation points (53), i.e., (m,n)=(m,0)(m,n){=}(m,0) or (m,n)=(0,n)(m,n){=}(0,n), the bifurcating branch is tangential to solutions of $D{\cal G}(0,\Lambda)[(\psi,0)]=0$ where $\cal G$} is the rhs of (51). We thus introduce the small parameter
@@
$\varepsilon u=\varepsilon A\psi_{m,n}+\varepsilon^{2}\sum_{(i,j)\in{\cal B}}A_{i,j}\psi_{i,j}+\mathrm{c.c.}+{\cal O}(\varepsilon^{3}),$ (59)
@@
Remark 4.3.
Near a bifurcation point from a stable branch, the critical mode determines the evolution, while all others are exponentially damped. Hence, in a standard AE setting, the signs of $a$, $\beta_{2}$ and $b$ in (61) determine the stability on the bifurcating branch. Here, from (61) we get stability information for steady states under area constraint, with $\lambda_{2}$ as extrinsic fixed parameter, but this can be extended to stability under area and volume constraints as follows. Close to the bifurcation point $\lambda_{2}(m,n)$ the solution reads $X(\varepsilon)={\cal C}_{L}-\varepsilon A\psi N$. Then the solution is linearly stable iff ${\partial}_{U}{\cal G}(X(\varepsilon),\Lambda)$ (which includes the $q_{1}$ area constraint) has only negative eigenvalues.
For admissible (area and volume preserving) perturbations $\phi:{\cal C}_{L}\rightarrow{\mathbb{R}}$ of $X(\varepsilon)$, we have the necessary and sufficient conditions
$\int_{X(\varepsilon)}H(\varepsilon A\psi)\phi\,{\rm d}S=0\quad\text{ and }\quad\int_{X(\varepsilon)}\varepsilon A\psi g^{-1/2}(\varepsilon A\psi)\phi\,{\rm d}S=0.$ (62)
For the mean curvature we have $H(\varepsilon A\psi)=\frac{1}{2}-\varepsilon A(\mu(m,n)+1)\psi+{\cal O}(\varepsilon^{2})$, and hence admissible perturbations must be orthogonal to $\psi$. In order to construct an admissible perturbation $\tilde{\phi}$, we take a Fourier mode $\psi_{kl}$ as a leading order term. Then we find that if ${\partial}_{u}{\cal A}(\varepsilon)\psi_{kl}={\cal O}(\varepsilon^{2})$ and ${\partial}_{u}{\cal V}(\varepsilon)\psi_{kl}={\cal O}(\varepsilon^{2})$, we can construct a $\phi=\psi_{kl}+g(\varepsilon)$, with $g={\cal O}(\varepsilon)$. Hence, for a branch described by $X(\varepsilon)={\cal C}_{L}-\varepsilon A\psi_{mn}N+{\cal O}(\varepsilon^{2})$, the only possible perturbations have a leading term with $(m,n)\neq(k,l)$. Thus, near bifurcation points associated with loss of stability of ${\cal C}_{L}$ we have the following results:
  • Branches bifurcating from a simple BP (pearling or wrinkling) are stable wrt (5).
  • Solutions on branches from a double BP (buckling and coiling) can only be destabilized by the other kernel vectors. We will see examples of this in §5. $\rfloor$
@@ 4.2.1 Pearling instability
    ${\hskip-11.38109pt}\frac{{\rm d}}{{\rm d}T}A=$ $\left(\frac{5k^{6}}{2}{+}\frac{\left(1{-}8c_{0}{-}3\lambda_{2}\right)k^{4}}{2}{+}\frac{\left(\frac{7}{2}{+}12c_{0}{+}\lambda_{2}\right)k^{2}}{2}{-}\frac{27}{4}{-}3\lambda_{2}\right)A{|A|}^{2}$
    $+\left(\frac{\left({-}4c_{0}{-}1\right)k^{2}}{2}{+}2\lambda_{2}{+}\frac{5}{2}\right)AA_{0}$
@@
$b=$ $\frac{5}{\left(k^{2}{-}1\right)k^{2}\left(4k^{4}{-}5k^{2}{+}4c_{0}{-}1\right)}\Bigg{(}\frac{3}{20}k^{14}{+}\left(4c_{0}{-}\frac{73}{20}\right)k^{12}{+}\left(\frac{8c_{0}}{5}{-}\frac{13}{10}\right)k^{10}$
${+}\left(\frac{32}{5}c_{0}^{2}{-}\frac{52}{5}c_{0}{+}\frac{73}{20}\right)k^{8}{+}\left(\frac{16}{5}c_{0}^{2}{-}8c_{0}{+}\frac{31}{5}\right)k^{6}{+}\left(\frac{4c_{0}}{5}{-}\frac{43}{20}\right)k^{4}{+}\left(\frac{12c_{0}}{5}{-}\frac{3}{2}\right)k^{2}\Bigg{)}.$
Recall that here $k=k_{m}=2m\pi/L$, hence $a<0$ for $L>2m\pi$, and similarly $b<0$ for non small $L$. For instance, $a\approx-0.6$ and $b\approx-3$ in Fig.5, where we compare the AE predictions for $m=1$ pearling with $c_{0}=0$ and $L=10$ with numerics. In any case for $ab<0$ we set $\beta_{2}=-1$ (subcritical case), and otherwise $\beta_{2}=1$ (supercritical case), giving the amplitude $A=\sqrt{\left|\frac{a}{b}\right|}$, with also $\alpha_{2}$ uniquely determined. The change of the volume can be approximated by inserting the amplitude Ansatz, giving $0=A_{0}+A^{2}+\beta_{2}$ at $\varepsilon^{2}\psi_{00}$, and comparison with (63) shows that preserving the area and the volume is not possible along the pearling branch.
@@ 4.2.2 Wrinkling instability
    $\frac{{\rm d}}{{\rm d}T}A=-6\left(\lambda_{2}-\frac{11}{4}\right)AA_{0}+\frac{1}{4}\left(-120\lambda_{2}+1170\right)\overline{A}A_{4}-3A\alpha_{2}-3\left(-\lambda_{2}+\frac{221}{4}\right){|A|}^{2}A$.
@@ 4.2.3 Coiling/buckling instability
For the instability with critical wave vector $(m,n)=(1,1)$ and (m,n)=(1,1)(m,n)=(1,-1) we choose the ansatz
@@
$\frac{{\rm d}}{{\rm d}T}A$ $=A(a\beta_{2}+b_{1}|A|^{2}+b_{2}|B|^{2}),$ (67a)
$\frac{{\rm d}}{{\rm d}T}B$ $=B(a\beta_{2}+b_{2}|A|^{2}+b_{1}|B|^{2}).$ (67b)
From that we can directly recover the two types of solutions: Coiling with $A=\sqrt{\left|\frac{a}{b_{1}}\right|}$ and $B=0$ or the other way round; buckling with $A=B=\pm\sqrt{\left|\frac{a}{b_{1}+b_{2}}\right|}$. Here again $\beta_{2}=\pm 1$, depending on the signs of $a$, $b_{1}$, and $b_{1}+b_{2}$.
We have $a=k_{m}^{2}$, but the computation of the coefficients $b_{1}$ and $b_{2}$ naturally is more cumbersome than for the scalar case (pearling and wrinkling), and therefore relegated to Appendix A. Here we instead briefly comment on the bifurcation directions, and stability of the bifurcating branches under volume and area preserving flows, see also Fig.3. The AEs are the same as in the classical $D_{4}$ symmetry problem [Hoy06, §4.3.1] and lead to the same direction of bifurcations. However the stability information gained is completely different, due to the constraints, cf. Remark 4.2. Namely, the Jacobian of the right hand side of (67) at a steady state $(A,B)$ is
@@
For $a>0$, on the coiling solution branch the eigenpairs are
$\sigma_{1}=(a\beta_{2}+3{\rm sign}(b_{1})|a\beta_{2}|)$, $v_{1}=(1,0)^{T}$, and $\sigma_{2}=(a\beta_{2}+\frac{b_{2}}{|b_{1}|}|a\beta_{2}|)$, $v_{2}=(0,1)^{T}$. (68)
The first direction $v_{1}$ is excluded by the constraints as it points along the coiling branch, and hence the coiling is stable for $\frac{b_{2}}{|b_{1}|}<-{\rm sign}(\beta_{2})$. For the buckling the Jacobian has the eigenpairs
$\sigma_{1}=2\frac{b_{2}+b_{1}}{|b_{1}+b_{2}|},v_{1}=(1,1)^{T}$, and $\sigma_{2}=\frac{b_{2}-b_{1}}{|b_{1}+b_{2}|},v_{2}=(1,-1)^{T}$. (69)
From the constraints again $v_{1}$ is excluded and hence the relevant eigenvalue is $\sigma_{2}$, which is negative for $b_{2}<b_{1}$. In §5 we give two comparisons of the predictions from (67) with numerics, namely $c_{0}=0$ and $L=10$ yielding $(a,b_{1},b_{2})\approx(0.39,-0.67,-0.82)$ in Fig.4, and $c_{0}=0.48$ and $L=10$ yielding $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$ in Fig.10. In both cases we get the correct branching and stability predictions, and also very good quantitative agreement.
@@
Figure 3: Schematic linearized stability diagram of the amplitude equation depending on $b_{1}$ and $b_{2}$. Orange stands for the coiling and green for the buckling branch. Thick lines indicate stable solutions.